OpenStudy (anonymous):

1/10k- 1/15=-1/6 solve for k

OpenStudy (texaschic101):

1/10k - 1/15 = -1/6 1/10k = -1/6 + 1/15 (common denominator is 30) 1/10k = - 5/30 + 2/30 1/10k = - 3/30 1/10k = - 1/10 -- divide by 1/10 k = 1/10k * - 10/1 k = - 10/10 k = -1 or you could just multiply the entire equation by the LCD, which is 30 1/10k - 1/15 = -1/6 --->(30) 3k - 2 = -5 -- add 2 to both sides 3k = -5 + 2 3k = -3 -- divide both sides by 3 k = -1